Evaluate the Reynold’s number at terminal velocity in air at 20oC, of a spherical dust particle of radius 1.0 X 10-5m and density 2.0 X 10-3Kgm-3.                           


1
Expert's answer
2021-06-29T13:03:06-0400

Gives

"\rho_{air}=1.2kg/m^3"

"\eta_{air}=1.8\times10^{-5}Nsec/m^2"

d="1.0\times10^{-5}m"

"\rho_{dust}=2.0\times10^{-3}kgm^{-3}"

"V_t=\frac{2r^2(\rho-\sigma)}{9\eta}g"

"V_t=\frac{2\times1\times10^{-5}\times(1.2-0.002)}{9\times1.8\times10^{-5}}\times9.8"

"V_t=1.44m\/sec"

Reynold's number

"R=\frac{\rho v_c d}{\eta}"

Put value

"R=\frac{1\times10^{-5} \times2.0\times10^{-3}\times1.44}{1.8\times10^{-5}}"

"R=1600"